% notes for the August 14th meeting. Purpose: Status of Deployment. Attendees: Jielun Sun, Britt Stephens, Steve Oncley, Sean Burns, Tony Delany, Joe, and Don Lenschow (at the very end). Some Items: ------------ * The ACD Balloon is up and collecting data (the group was out in the field so no one was there to report the status). * Jielun, Britt, Alex, Andrew, ATD, and Sean were out in the field yesterday (aug13). Some new photos are on the web page. * Hydra is almost ready for deployment. Tim is leaving this weekend. Tim has all the ground lines laid out and needs to get the vertical lines put out. There are only a few more hours left for the assembly. Pierre (from ACD) is supposed to help Tony get the LiCors hooked up. The hydra is supposed to be deployed next week. * Tram is to be setup at FL next week sometime for a test run before deployment. It is expected to be deployed by the end of the month. * The cart Britt was planning to use to transport the cal gas cylinders will need to be redesigned using smaller cylinders due to rough terrain on the trails. Britt will be out of town until Thur of next week (Aug 22). He plans to go up on Friday (Aug 23) to the field site. * The thermocouples are installed on the CU tower, but a cable to get power to the data logger and transmit the data is still needed. Steve O. was going to help Sean get this. * The ATD towers are up with instrumentation. * Andrew collected the differential GPS data and they are now available on the webpage. * Jielun, Britt, and Sean drove up the road to investigate a light-colored area on the IR image from LTER. This turned out to be the swampy area which had low shrubs and was above Como Creek...some photos of the area were taken. Some info about this will eventually be added to the webpage.